OpenStudy (anonymous):

What are the x- and y- intercepts of the following equation: y = 56 – x

OpenStudy (shayaan_mustafa):

Hi mccutchen :) How are you dear?

OpenStudy (shayaan_mustafa):

For x-intercept: Put y=0 y=56-x 0=56-x x=56 For y-intercept: Put x=0 y=56-x y=56-0 y=56

OpenStudy (shayaan_mustafa):

Final answer is, x-intercept (56,0) y-intercept (0,56)

OpenStudy (anonymous):

Thank you! :)

OpenStudy (shayaan_mustafa):

Welcome.